### v1

@@ 4 Improving Training-free Guidance
This section presents enhancement techniques for improving training-free guidance, with theoretical justifications and empirical evidence.
@@
     $\bm{x}_{t-1}=\bm{x}_{t-1}-\eta\cdot\frac{\bm{g}_{t}}{\|\bm{g}_{t}\|_{2}}$
@@ 4.1 Randomized Augmentation
As established by Proposition 2, the introduction of Gaussian perturbations enhances the Lipschitz property of a neural network. A direct application of this principle involves creating multiple noisy instances of an estimated clean image and incorporating them into the guidance network, a method analogous to the one described in (6). However, given the high-dimensional nature of image data, achieving a satisfactory approximation of the expected value necessitates an impractically large number of noisy copies. To circumvent this issue, we propose an alternative strategy that employs a diverse set of data augmentations in place of solely adding Gaussian noise. This approach effectively introduces perturbations within a lower-dimensional latent space, thus requiring fewer samples. The suite of data augmentations utilized, denoted by $\mathcal{T}$, is derived from the differentiable data augmentation techniques outlined in [46], which encompasses transformations such as translation, resizing, color adjustments, and cutout operations. The details are shown in Algorithm 1 and the rationale is shown in the following proposition.
Proposition 4
@@
(Random augmentation improves smoothness) Given a bounded non-Lipschitz loss function $\ell(\bm{x})$, the loss ^$\hat{\ell}(\bm{x})=\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\ell(\bm{x}+\bm{\epsilon})]$ is
@@
The proof is shown in Appendix 8.5. Echoing the experimental methodology delineated in Section 3.2, we present an analysis of the accumulated gradient effects when applying random augmentation to a ResNet-50 model. Specifically, we utilize a set of $|\mathcal{T}|=10$ diverse transformations as our augmentation strategy. The results of this experiment are visualized in Figure 1(d), where the target object’s color and shape emerges in the gradient profile. This observation suggests that the implementation of random augmentation can alleviate the adversarial gradient issue.
4.2 Adaptive Gradient Scheduling
In Section 3.1, we established that the training guidance process can be conceptualized as the optimization of the loss function corresponding to the guidance network. Given this interpretation, it stands to reason that adopting a more sophisticated optimizer could accelerate the convergence of the guidance. Notably, the fundamental operation in guidance involves calculating the gradient of the guidance network with respect to the input, a procedure reminiscent of white-box adversarial attacks. Drawing inspiration from this parallel, we adopt the well-known Projected Gradient Descent (PGD) optimization technique—commonly employed in crafting adversarial examples—as our optimizer of choice, which is shown in Algorithm 2. We implement PGD within the context of a training-free guidance framework called FreeDoM [44] and benchmark the performance of this implementation using the DDIM sampler with $50$ steps, ensuring identical seed initialization across the different methods for consistency.
As shown in Figure 3, FreeDoM is unable to effectively guide the generation process when faced with a significant discrepancy between the unconditional generation and the specified condition. An illustrative example is the difficulty in guiding the model to generate faces oriented to the left when the unconditionally generated faces predominantly orient to the right, as shown in the middle column of the figure. Conversely, when the unconditionally generated faces orient forward, the condition for any face orientation is more easily fulfilled. This challenge, which arises due to the insufficiency of $50$ steps for convergence under the condition, is ameliorated by substituting gradient descent with PGD, thereby illustrating the benefits of employing an advanced optimizer in the guidance process.
Figure 3: FreeDoM [44] with and without PGD trick. DDIM-$50$ is used for sampling.
4.3 Resampling Trick
Algorithm 3 Resampling
@@ 4.3 Resampling Trick
The technique of resampling, also referred to as “time travel”, has been proposed as a solution to complex generative problems [24], facilitating successful training-free guidance in tasks such as CLIP-guided ImageNet generation and layout guidance for stable diffusion, as illustrated in Figure 2 of [44] and Figure 8 of [2], respectively. The procedure of the resampling trick is shown in Algorithm 3, which involves recursive execution of individual sampling steps.
In this subsection, we endeavor to analyze the role of resampling in training-free guidance and thus complete the overall picture. During the diffusion process, samples can diverge from their target distributions due to various types of sampling error, such as adversarial gradient and discretization error. A critical component of the resampling process entails the reintroduction of random noise to transition the sample $\bm{x}_{t-1}$ back to $\bm{x}_{t}$, effectively resetting the sampling state. The forthcoming lemma substantiates the premise that the act of adding identical noise to two vectors serves to diminish the disparity between the samples.
Lemma 1
[…]
Following Lemma 1, the next proposition demonstrates that the total variation distance between the sampled distribution $q_{t}(\bm{x}_{t}^{i})$, as obtained in the $i$-th resampling of Algorithm 3, and the ground-truth distribution $p_{t}(\bm{x}_{t})$, is progressively reduced through recursive application of the resampling steps.
Proposition 5
@@
[…]
where $h_{\max}$ is defined in Proposition 3.

### v2

@@ 4 Improving Training-free Guidance
In this section, we propose to adopt random augmentation to mitigate the adversarial gradient issue, and Polyak step size to mitigate the convergence issue. Time travel was adopted in previous studies [2, 46, 14] without theoretical justifications. We theoretically analyze this trick in Appendix C.2, which completes the picture for the tricks to improve training-free guidance.
@@
     $\bm{x}_{t-1}=\bm{x}_{t-1}-\eta\cdot\frac{\|\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\|}{\|\bm{g}_{t}\|_{2}^{2}}\cdot\bm{g}_{t}$
@@ 4.1 Random Augmentation
As established by Proposition 3.2, the introduction of Gaussian perturbations enhances the Lipschitz property of a neural network. A direct application of this principle involves creating multiple noisy instances of an estimated clean image and incorporating them into the guidance network, a method analogous to the one described in (6). However, given the high-dimensional nature of image data, achieving a satisfactory approximation of the expected value necessitates an impractically large number of noisy copies. To circumvent this issue, we propose an alternative strategy that employs a diverse set of data augmentations in place of solely adding Gaussian noise. This approach effectively introduces perturbations within a lower-dimensional latent space, thus requiring fewer samples. The suite of data augmentations utilized, denoted by $\mathcal{T}$, is derived from the differentiable data augmentation techniques outlined in [48], which encompasses transformations such as translation, resizing, color adjustments, and cutout operations. The details are shown in Algorithm 1 and the rationale is shown in the following proposition.
Proposition 4.1.
@@
(Random augmentation improves smoothness) Given a bounded non-Lipschitz loss function $\ell(\bm{x})$, the loss ^$\hat{\ell}(\bm{x})=\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\ell(\bm{x}+\bm{\epsilon})]$ is $C\int_{\mathbb{R}^{n}}\|\nabla p(\bm{t})\|_{2}{\rm d}\bm{t}$-Lipschitz and its gradient is $C\int_{\mathbb{R}^{n}}\|\nabla^{2}p(\bm{t})\|_{\text{op}}{\rm d}\bm{t}$-Lipschitz.
The proof is shown in Appendix D.5. Echoing the experimental methodology delineated in Section 3.2, we present an analysis of the accumulated gradient effects when applying random augmentation to a ResNet-50 model. Specifically, we utilize a set of $|\mathcal{T}|=10$ diverse transformations as our augmentation strategy. The results of this experiment are visualized in Figure 2(d), where the target object’s color and shape emerges in the gradient profile. This observation suggests that the implementation of random augmentation can alleviate the adversarial gradient issue. The efficiency of random augmentation is further discussed in Appendix C.3.
4.2 Polyak Step Size
(a) When the initialization is proximate to the specified conditions, both step sizes perform satisfactorily.
(b) When the initialization deviates from the conditions, only the Polyak step size guides effectively.
Figure 3: The effects of step size.
In Section 3.1, we analyzed training-free guidance from the optimization perspective. To accelerate the convergence, gradient step size should be adaptive to the gradient landscape. We adopt Polyak step size, which has near-optimal convergence rates under various conditions [13]. The algorithm is shown in Algorithm 2 and the term $\|\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\|$ is used to both estimate the gap to optimal values and balance the magnitude of diffusion term and guidance term.
We implement Polyak step size within the context of a training-free guidance framework called FreeDoM [46] and benchmark the performance of this implementation using the DDIM sampler with $50$ steps. As shown in Figure 3, FreeDoM is unable to effectively guide the generation process when faced with a significant discrepancy between the unconditional generation and the specified condition. An illustrative example is the difficulty in guiding the model to generate faces oriented to the left when the unconditionally generated faces predominantly orient to the right, as shown in Figure 3(b). This challenge, which arises due to the insufficiency of $50$ steps for convergence under the condition, is ameliorated by substituting gradient descent with adaptive step size, thereby illustrating the benefits of employing a better step size in the guidance process.
5 Experiments
In this section, we evaluate the efficacy of our proposed techniques across various diffusion models and guidance conditions. We compare our methods with established baselines: Universal Guidance (UG) [2], Loss-Guided Diffusion with Monte Carlo (LGD-MC) [35], Training-Free Energy-Guided Diffusion Models (FreeDoM) [46], and Manifold Preserving Guided Diffusion (MPGD) [14]. LGD-MC utilizes (6) while UG and FreeDoM are built on (5). MPGD utilizes an auto-encoder to ensure the manifold constraints. Furthermore, time travel trick (Algorithm 3) is adopted in UG, FreeDoM, and MPGD to improve the performance. Please refer to Appendix B for details. For the sampling method, DDIM with $100$ steps is adopted as in [46, 35]. The method “Ours” is built on FreeDoM, with Polyak step size and random augmentation.
5.1 Guidance to CelebA-HQ Diffusion
We randomly select $1000$ samples each of segmentation maps, sketches, and text descriptions. The comparative results are presented in Table 1. Consistent with [46], the time-travel number for all methods is set to $s=1$. Figure 4 displays a random selection of the generated images. More image samples are provided in the supplementary materials. We find that the baselines failed to guide if the condition differs from unconditionally generated images significantly, as discussed in Section 4.2.
Methods Segmentation maps Sketches Texts
Distance↓ FID↓ Distance↓ FID↓ Distance↓ FID↓
UG [2] 2247.2 39.91 52.15 47.20 12.08 44.27
LGD-MC [35] 2088.5 38.99 49.46 54.47 11.84 41.74
FreeDoM [46] 1657.0 38.65 34.21 52.18 11.17 46.13
MPGD-Z [14] 1976.0 39.81 37.23 54.18 10.78 42.45
Ours 1575.7 33.31 30.41 41.26 10.72 41.25
5.2 Guidance to ImageNet Diffusion
Methods LGD-MC [35] UG [2] FreeDoM [46] MPGD-Z [14] Ours
CLIP Score↑ 24.3 25.7 25.9 25.1 27.7
For the unconditional ImageNet diffusion, we employ text guidance in line with the approach used in FreeDoM and UG [2, 46]. We utilize CLIP-B/16 as the image and text encoder, with cosine similarity serving as the loss function to measure the congruence between the image and text embeddings. To evaluate performance and mitigate the potential for high-scoring adversarial images, we use CLIP-L/14 for computing the CLIP score. In FreeDoM and MPGD-Z, resampling is conducted for time steps ranging from $800$ to $300$, with the time-travel number fixed at $10$, as described in [46]. Given that UG resamples at every step, we adjust its time-travel number $s=5$ to align the execution time with that of FreeDoM. The textual prompts for our experiments are sourced from [23]. The comparison of different methods is depicted in Table 2. The corresponding randomly selected images are illustrated in Figure 5. The table indicates that our method achieves the highest consistency with the provided prompts. As shown in Figure 5, LGD-MC and MPGD tend to overlook elements of the prompts. Both UG and FreeDoM occasionally produce poorly shaped objects, likely influenced by adversarial gradients. Our approach addresses this issue through the implementation of random augmentation. Additionally, none of the methods successfully generate images that accurately adhere to positional prompts such as “left to” or “below”. This limitation is inherent to CLIP and extends to all text-to-image generative models [39]. More image samples are provided in the supplementary materials.
5.3 Guidance to Human Motion Diffusion
Methods “Backwards” “Balanced Beam” “Walking” “Jogging”
Loss↓ CLIP↑ Loss↓ CLIP↑ Loss↓ CLIP↑ Loss↓ CLIP↑
In this subsection, we extend our evaluation to human motion generation using the Motion Diffusion Model (MDM) [38], which represents motion through a sequence of joint coordinates and is trained on a large corpus of text-motion pairs with classifier-free guidance. We apply the targeting guidance and object avoidance guidance as described in [35]. Let $\bm{x}_{0}(t)$ denote the joint coordinates at time $t$, $\bm{y}_{t}$ the target location, $\bm{y}_{\text{obs}}$ the obstacle location, $r$ the radius of the objects, and $T$ the total number of frames. The loss function is defined as follows:
Our experimental configuration adheres to the guidelines set forth in [35]. We assess the methods using the targeting loss (the first term in (8)), the object avoidance loss (the second term in (8)), and the CLIP score calculated by MotionCLIP [37]. In this application, MPGD-Z cannot be applied as there are no auto-encoder. MPGD w/o proj suffers from the shortcut and cannot achieve good performance, as discussed in Appendix B.2. In our method, random augmentation is omitted because the guidance is not computed by neural networks so the adversarial gradient issues are not obvious. The quantitative results of our investigation are summarized in Table 3, while Figure 6 showcases randomly selected samples. Our methods exhibit enhanced control quality over the generated motion. The videos are provided in the supplementary materials.
6 Conclusions
In this paper, we conducted a comprehensive investigation into training-free guidance, which employs pre-trained diffusion models and guides them using the off-the-shelf trained on clean images. Our exploration delved into the underlying mechanisms and fundamental limits of these models. Moreover, we proposed a set of enhancement techniques and verified their effectiveness both theoretically and empirically. Limitations. Despite our efforts to mitigate the shortcomings of training-free methods and enhance their performance, certain limitations remain. Notably, the refined training-free guidance still necessitates a higher number of NFEs when compared with extensive training methods such as classifier-free guidance. This is because adversarial gradient cannot be fully eliminated without training. Ethical Consideration. Similar to other models designed for image creation, our model also has the unfortunate potential to be used for creating deceitful or damaging material. We pledge to restrict the usage of our model exclusively to the realm of research to prevent such misuse.
Appendix A Related Works
A.1 Training-Free Control for Diffusion
The current training-free control strategies for diffusion models can be divided into two primary categories. The first category is the investigated topic in this paper, which is universally applicable to universal control formats and diffusion models. These methods predict a clean image, subsequently leveraging pre-trained networks to guide the diffusion process. Central to this approach are the algorithms based on (5), which have been augmented through techniques like time-travel [2, 46] and the introduction of Gaussian noise [35]. Extensions of these algorithms have found utility in domains with constrained data-condition pairs, such as molecule generation [12], and in scenarios necessitating zero-shot guidance, like open-ended goals in offline reinforcement learning [42]. In molecular generation and offline reinforcement learning, they outperform training-based alternatives as additional training presents challenges. This paper delves deeper into the mechanics of this paradigm and introduces a suite of enhancements to bolster its performance. The efficacy of our proposed modifications is demonstrated across image and motion generation, with promising potential for generalization to molecular modeling and reinforcement learning tasks.
The second category of training-free control is tailored to text-to-image or text-to-video diffusion models, which is based on insights into their internal backbone architecture. For instance, object layout and shape have been linked to the cross-attention mechanisms [15], while network activations have been shown to preserve object appearance [40]. These understandings facilitate targeted editing of object layout and appearance (Diffusion Self-Guidance [10]) and enable the imposition of conditions in ControlNet through training-free means (FreeControl [29]). Analyzing these methodologies is challenging due to their reliance on emergent representations during training. Nonetheless, certain principles from this paper remain relevant; for example, as noted in Proposition 3.3, these methods often necessitate extensive diffusion steps, with instances such as [29, 10] employing $1000$ steps. A thorough examination and refinement of these techniques remain an avenue for future research.
A.2 Training-based Gradient Guidance
The training-based gradient guidance paradigm, such as classifier guidance, is a predominant approach for diffusion guidance. The core objective is to train a time-dependent network that approximates $p_{t}(\bm{y}|\bm{x}_{t})$ in the RHS of (2), and to utilize the resulting gradient as guidance. The most well-known example is classifier guidance, which involves training a classifier on noisy images. However, classifier guidance is limited to class conditions and is not adaptable to other forms of control, such as image and text guidance. To address this limitation, there are two main paradigms. The first involves training a time-dependent network that aligns features extracted from both clean and noisy images, as described by [22]. The training process is outlined as follows:
$\min_{\psi}\quad\mathbb{E}_{p(\bm{x}_{0},\bm{x}_{t})}d(f_{\psi}(\bm{x}_{t},t),f_{\phi}(\bm{x}_{0})),$
where $d(\cdot,\cdot)$ represents a loss function, such as cross-entropy or the $\ell_{2}$ norm. If time-dependent networks for clean images are already available, training can proceed in a self-supervised fashion without the need for labeled data. The second paradigm, as outlined by [19], involves training an energy-based model to approximate $p_{t}(\bm{y}|\bm{x}_{t})$. The training process is described as follows:
$\min_{\psi}\quad\mathbb{E}_{p(\bm{x}_{0},\bm{x}_{t})}|\ell(f_{\psi}(\bm{x}_{t},t),\bm{y})-\ell(f_{\phi}(\bm{x}_{0}),\bm{y})|.$
However, it is observed in [25] that none of these methods can accurately approximate the true energy in (4). The authors of [25] propose an algorithm to learn the true energy. The loss function is a contrastive loss
$\min_{\psi}\quad\mathbb{E}_{p(\bm{x}_{0}^{i},\bm{x}_{t}^{i})}\exp(-\ell(f_{\phi}(\bm{x}_{0}),\bm{y}))\left[-\sum_{i=1}^{K}\log\frac{\exp(\ell(f_{\psi}(\bm{x}_{t}^{i},t),\bm{y}^{i})}{\sum_{j=1}^{K}\exp(-\ell(f_{\psi}(\bm{x}_{t}^{i},t),\bm{y}^{i}))}\right],$
where $(\bm{x}_{0}^{i},\bm{x}_{t}^{i})$ are $K$ paired data samples from $p(\bm{x}_{0}^{i},\bm{y}^{i})$. Theorem 3.2 in [25] proves that the optimal $f_{\psi^{*}}$ satisfied that $\nabla_{\bm{x}_{t}}\ell(f_{\psi^{*}}(\bm{x}_{t}^{i},t),\bm{y}^{i})=\nabla_{\bm{x}_{t}}p_{t}(\bm{y}|\bm{x}_{t})$.
Although this paper focuses on training-free guidance, the findings in this paper can be naturally extended to all training-based gradient guidance schemes. Firstly, the issue of adversarial gradients cannot be resolved without additional training; hence, all the aforementioned methods are subject to adversarial gradients. Empirical evidence for this is presented in Fig. 2, which illustrates that the gradients from an adversarially robust classifier are markedly more vivid than those from time-dependent classifiers. Consequently, it is anticipated that incorporating additional adversarial training into these methods would enhance the quality of the generated samples. Secondly, since these methods are dependent on gradients, employing a more sophisticated gradient solver could further improve their NFEs.
A.3 Adversarial Attack and Robustness
Adversarial attacks and robustness constitute a fundamental topic in deep learning [4]. An adversarial attack introduces minimal, yet strategically calculated, changes to the original data that are often imperceptible to humans, leading models to make incorrect predictions. The most common attacks are gradient-based, for example, the Fast Gradient Sign Method (FGSM) [11], Projected Gradient Descent (PGD) [27], Smoothed Gradient Attacks [1], and Momentum-Based Attacks [9]. An attack is akin to classifier guidance or training-free guidance, which uses the gradient of a pre-trained network for guidance. Should the gradient be adversarial, the guidance will be compromised. This paper establishes the relationship between training-free loss-guided diffusion models and adversarial attacks in two ways. Firstly, we prove that training-free guidance is more sensitive to an adversarial gradient. Secondly, in Section 4.2, we demonstrate that borrowing an adaptive gradient scheduler can improve convergence. The optimizers from adversarial attack literature may also expedite the convergence of the diffusion ODE.
Appendix B Baselines and Experimental Settings
B.1 Details of the Baselines
We use the following training-free diffusion guidance methods as baselines for comparison:
  • Universal Guidance (UG) [2] employs guidance as delineated in (5) and uses time-travel strategies outlined in Algorithm 3 to enhance performance. The time-travel trick is used for all time steps $t$.
  • FreeDoM [46] is also founded on (5) and time-travel trick. In addition, FreeDoM incorporates a time-dependent step size for each gradient guidance and judiciously selects the diffusion step for executing time-travel trick.
  • Loss-guided Diffusion with Monte Corlo (LGD-MC) [35] utilizes guidance from (6) and we set $n=10$ in the experiments.
  • Manifold Guided Preserving Diffusion (MPGD) [14] takes the derivative with respect to estimated clean image $\mathbb{E}[\bm{x}_{0}|\bm{x}_{t}]$ instead of $\bm{x}_{t}$. Let $\bm{x}_{0|t}=\mathbb{E}[\bm{x}_{0}|\bm{x}_{t}]$, MPGD steps are expressed as following:
    $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t}):=-\sqrt{\alpha_{t-1}}\nabla_{\bm{x}_{0|t}}\log\left[\exp(-\ell(f_{\phi}(\bm{x}_{0|t}),\bm{y})\right].$
    MPGD-Z adopts an additional auto-encoder to preserve manifold constraints. The details procedures of MPGD-Z are described in Algorithm 3 of [14].
B.2 MPGD for Motion Diffusion
For the process of motion diffusion, the application of both MPGD-Z and MPGD-AE is precluded due to the absence of pretrained auto-encoders specific to motion diffusion. An implementation of MPGD without projection (MPGD w/o proj) was attempted for motion diffusion; however, it was unsuccessful in accurately navigating towards the target. This failure is attributed to the presence of spurious correlations within the targeting loss specific to MPGD, a phenomenon not observed in the other baseline methodologies. Sepcifically, the gradient formulation in MPGD is detailed as follows:
@@ B.2 MPGD for Motion Diffusion
The gradient in FreeDoM and other methods is given by
[…]
Analysis of the aforementioned equations reveals that, within the MPGD framework, only the final motion step is influenced by the gradient, a characteristic not shared by alternative methodologies. Consequently, this exclusive focus on the last step results in disproportionately strong guidance at this juncture, while earlier steps suffer from a lack of directional input. This imbalance may adversely affect the overall quality of the samples produced. Empirical observations substantiate that MPGD struggles to achieve targeted outcomes when a nuanced adjustment of step size is required. Given these limitations, MPGD has been excluded from the comparative analysis presented in Table 3.
B.3 Prompts for Motion Diffusion
We follow the prompts and evaluation settings in [35]. The prompts are (i) “the person is walking backwards”; (ii) “a person walking around in a balance beam”; (iii) “the person is walking”; (iv) “the person is jogging”. We consider three different directions for each prompt, and each direction has $10$ random seeds, the metrics are then averaged together over the $30$ synthesized motions.
Appendix C More Discussions
C.1 Concentration of Estimated Clean Samples
It has been demonstrated in [7] that, given a fixed $\bm{x}_{0}$, the distribution of the noisy data $\bm{x}_{t}$ is concentrated on a spherical shell. Furthermore, an extension of this theorem presented in the ’High Dimensional Probability’ textbook by Vershynin [41] elucidates that the conditional distribution $q(\bm{x}_{0}|\bm{x}_{t})$ also exhibits concentration on a spherical shell, provided that its coordinates are sub-Gaussian.
Theorem C.1.
(Theorem 3.1.1 in [41]) Denote $\bm{x}_{0}=\bm{x}_{t}+\bm{g}\in\mathbb{R}^{n}$, where $g=(g_{1},\cdots,g_{n})$. Assume that $g_{i}$ are independent identically distributed, $\mathbb{E}[g_{i}^{2}]=\frac{r_{t}^{2}}{n}$ and there exists constant $c_{1}$ such that $\mathbb{P}[g_{i}\geq t]\leq\exp(-c_{1}t^{2})$, then we have
$\mathbb{P}[|\|\bm{x}_{t}-\bm{x}_{0}\|_{2}^{2}-r_{t}^{2}|\geq t]\leq\exp(-c_{2}nt^{2}),$
where $c_{2}$ is a constant.
Theorem C.1 establishes that the distribution $q(\bm{x}_{0}|\bm{x}_{t})$ exhibits concentration on a spherical shell at an exponential rate. For high-dimensional data, such as images, it is reasonable to infer that $\bm{x}_{0}$ is predominantly situated on this spherical shell.
C.2 Time-travel Trick
Algorithm 3 Time Travel
@@ C.2 Time-travel Trick
The technique of time-travel, also referred to as “resampling”, has been proposed as a solution to complex generative problems [26], facilitating successful training-free guidance in tasks such as CLIP-guided ImageNet generation and layout guidance for stable diffusion, as illustrated in Figure 2 of [46] and Figure 8 of [2], respectively. The procedure of the time-travel trick is shown in Algorithm 3, which involves recursive execution of individual sampling steps.
In this subsection, we endeavor to analyze the role of time-travel trick in training-free guidance and thus complete the overall picture. During the diffusion process, samples can diverge from their target distributions due to various types of sampling error, such as adversarial gradient and discretization error. A critical component of the time-travel trick process entails the reintroduction of random noise to transition the sample $\bm{x}_{t-1}$ back to $\bm{x}_{t}$, effectively resetting the sampling state. The forthcoming lemma substantiates the premise that the act of adding identical noise to two vectors serves to diminish the disparity between the samples.
Lemma C.2.
(Adding noise removes the discrepancy; Lemma 7 of [44]) Suppose $\bm{n}_{x},\bm{n}_{y}\sim\mathcal{N}(0,\sigma^{2}\bm{I})$, and $\bm{x}^{\prime}=\bm{x}+\bm{n}_{x},\bm{y}^{\prime}=\bm{y}+\bm{n}_{y}$. Then we have $\text{TV}(\bm{x}^{\prime},\bm{y}^{\prime})=1-2Q(r)$, where $Q(\cdot)$ is Gaussian tail probability defined in Definition D.3 in appendix and $r=\|\bm{x}-\bm{y}\|/(2\sigma)$.
Following Lemma C.2, the next proposition demonstrates that the total variation distance between the sampled distribution $q_{t}(\bm{x}_{t}^{i})$, as obtained in the $i$-th time travel of Algorithm 3, and the ground-truth distribution $p_{t}(\bm{x}_{t})$, is progressively reduced through recursive application of the time-travel steps.
Proposition C.3.
@@
$\text{TV}(q_{t}(\bm{x}_{t}^{s}),p_{t}(\bm{x}_{t}))\leq\left(1-2Q\left(\frac{\sqrt{\beta_{t}}B}{2\sqrt{1-\beta_{t}}}\right)\right)^{s-1}\text{TV}(q_{t}(\bm{x}_{t}^{1}),p_{t}(\bm{x}_{t}))+O(h_{\max}),$
where $h_{\max}$ is defined in Proposition 3.3.